# ITCMA: A Generative Agent Based on a Computational Consciousness Structure

### v1

@@ B.1. Alfworld
B.2. Quadruped robots in the real world
@@
Appendix C Explanation for High-Level Functions of Consciousness
@@
C.1. Self-Awareness
In the field of developmental psychology, a mirror recognition task is sometimes claimed to be a decisive test of self-awareness. It has been claimed that self-awareness emerges only when a child (about 18 months old) is able to recognize his or herself in a mirror [71]. Carruthers believed that animals and children younger than 18 months do not have self-awareness, as they lack the dimension of subjectivity. In his view, they cannot see the existence of their own mental states, nor is there any “feeling” that makes them aware of sensations of pain or pleasure [72; 73]. He also found that although some organisms only have the weakest concept of themselves as subjects with continuous thinking and experience, they may still have a conscious mental state. Thus, he believes that it is quite effective to talk about the conscious mental state of organisms that lack self-awareness.

### v2

@@ C.1. Alfworld
C.2. Quadruped robots in the real world
@@
Appendix D Explanation for High-Level Functions of Consciousness
@@
D.1. Self-Awareness
In the field of developmental psychology, a mirror recognition task is sometimes claimed to be a decisive test of self-awareness. It has been claimed that self-awareness emerges only when a child (about 18 months old) is able to recognize his or herself in a mirror [73]. Carruthers believed that animals and children younger than 18 months do not have self-awareness, as they lack the dimension of subjectivity. In his view, they cannot see the existence of their own mental states, nor is there any “feeling” that makes them aware of sensations of pain or pleasure [74; 75]. He also found that although some organisms only have the weakest concept of themselves as subjects with continuous thinking and experience, they may still have a conscious mental state. Thus, he believes that it is quite effective to talk about the conscious mental state of organisms that lack self-awareness.